Your firm needs a computerized machine tool lathe which costs $97,752 and requires $8,850 in maintenance expense for each year of its 5-year life. After five years, this machine will be replaced. The machine falls into the MACRS (Links to an external site.) 5-year class life category. Assume a tax rate of 25 percent and a discount rate of 12 percent. If the lathe can be sold for $11,070 at the end of year 5, what is the after-tax salvage value? Below are the MACRS rates for the 5-year recovery period. DO NOT USE DOLLAR SIGNS OR COMMAS IN YOUR ANSWER. ENTER YOUR ANSWER TO THE NEAREST DOLLAR (e.g. 1250).

Year MACRS Rate
1 20.00%
2 32.00%
3 19.20%
4 11.52%
5 11.52%
6 5.76%
Answer #1

rate positively ..

Computation of after tax salvage value -
a Book value after 5 year =           5,631
97,752*5.76%
b Sales price =        11,070
c=b-a Profit on sale =           5,439
d=c*25% Tax on profit =           1,360
e=b-d Post tax salvage value =           9,710
Ans =           9,710